OpenStudy (anonymous):

Maxima and minina... Find the positive numbers whose sum is 110 and whose product is a maximum.

OpenStudy (anonymous):

55,55

OpenStudy (anonymous):

55 $ 55

OpenStudy (anonymous):

55 and 55.... use :x+y=110 next let us say p=xy for product to be max dp/dx should be -ve and here p=x(110-x) u'll see that x=55 and y=55

OpenStudy (lalaly):

Let one number be called x. The other is 110-x. f(x) = x(110-x) f(x) = 110x - x^2 Now take the derivative f '(x) =110 - 2x make it =0 110- 2x = 0 110= 2x x = 110/2 so the other number is 110- 110/2 = 110/2. The two numbers are 110/2 and 110/2. =55 and 55